# Soft decision trees for survival analysis

### v1

@@ 3.3 Parametric and spline-based semiparametric models for survival functions
In statistical modeling, several alternatives have been proposed for the estimation of the survival and hazard functions. Traditionally, medical researchers and biostatisticians have relied on Cox regression (Cox, 1972) to handle censored survival data. This model assumes that the hazard rate for an event is a linear combination of the features, with the hazard function for an input vector $\mathbf{x}$ expressed as:
@@
In the Cox model, the choice of $h_{0}(t)$ is critical since its estimation in high-dimension is prone to overfitting and may complicate the validation (Gelfand et al., 2000; Royston and Parmar, 2002). Considering more sophisticated hazard functions can lead to more accurate estimates and offer deeper insights into the phenomenon under investigation (e.g., capturing the progression of a disease over time in medical studies). Moreover, the classical proportional hazards assumption of the Cox model may not be suitable for all contexts [^2] Although the Cox model allows for non-proportional hazard assumption, such as through time-varying regression coefficients, there is no widely accepted or intuitive method for implementing this adjustment..
To overcome these limitations, more flexible statistical approaches, including parametric distribution and spline-based semiparametric ones, have been proposed. These approaches are well-suited to capture complex patterns in the data, such as time-varying hazards, interval censoring, frailties, and multiple responses across diverse datasets or time scales (Aalen, 2008).
@@ 3.3.1 Parametric distribution models
All parameters may be linearly dependent on the features through a transformation $g$ such that $g(\alpha_{r}(\mathbf{x}))=\gamma_{0}+\bm{\gamma}^{T}\mathbf{x}$ for $r=1,\dots,R$. Typically, $g$ is the logarithm function for positive parameters, and the identity function for parameters which are unrestricted in sign. For a review of the parametric distributions, the reader is referred to (Lawless, 2014; Klein et al., 2014).
In this work, we assume that only the primary parameter $\mu$ linearly depends on the features (i.e., $\text{log}(\mu(\mathbf{x}))=\gamma_{0}+\bm{\gamma}^{T}\mathbf{x}$). Thus, for each leaf node $n\in\tau_{L}$, the vector $\bm{\beta}_{n}$ contains $p+1+m$ components. The first $p+1$ components correspond to the number of variables related to $\mu$, and the $m=R-1$ remaining ones to the ancillary parameters.
@@ 3.3.2 Spline-based semiparametric models
For the estimation of the semiparametric survival functions, we adopt the methodology proposed by Royston and Parmar (2002), which assumes flexible spline-based models to take into account proportional hazards and proportional odds. Other options (see, e.g., Younes and Lachin (1997); Shen (1998); Royston and Parmar (2011); Bremhorst and Lambert (2016)) that preserve the smoothness of formulation (2) can also be considered.
Given any input vector $\mathbf{x}$, Royston and Parmar (2002) assume that the transformation of the survival function $g(S_{\mathbf{x}}(t;\bm{\beta}))$ is equal to the transformation of a baseline survival function $S_{0}(t)$ plus a linear combination of the features:
$g(S_{\mathbf{x}}(t;\bm{\beta}))=g(S_{0}(t))+\bm{\gamma}^{T}\mathbf{x},$
@@
As in Aranda-Ordaz (1981), we consider:
$g(x;\theta)=\text{log}\frac{x^{-\theta}-1}{\theta}$
and we derive the proportional odds when $\theta=1$, and the proportional hazards when $\theta\rightarrow 0$. As in Royston and Parmar (2002), we estimate the baseline survival function $S_{0}(t)$, by modeling the logarithm of the baseline cumulative odds or hazard function as a natural cubic spline of logarithmic time $y=\text{log}(t)$.
This leads to the Proportional Odds (PO) spline model:
@@
The complexity of such splines is determined by the number of degrees of freedom, which is equal to $m+1$. For the placement of internal knots, Royston and Parmar (2002) propose to use the percentiles of the distribution of uncensored log survival times. Further details can be found in Table 1 of Royston and Parmar (2002).
For any leaf node $n\in\tau_{L}$ and any right-censored data point $(\textbf{x},t,c)$, the corresponding negative log likelihood $L^{-}_{n}$ under the PO assumption is:
$L^{-}_{n}({\mathbf{x}},t,c;\bm{\beta}_{n})=\begin{cases}-log\left[\frac{1}{t}\frac{\text{d}s(y;\bm{\eta}_{n})}{\text{d}y}\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x})(1+\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))^{-2}\right]\hskip 14.22636pt\text{if }c=1\\ -log\left[(1+\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))^{-1}\right]\hskip 156.49014pt\text{if }c=0,\end{cases}$
whereas under the PH assumption it is as follows:
$L^{-}_{n}({\mathbf{x}},t,c;\bm{\beta}_{n})=\begin{cases}-log\left[\frac{1}{t}\frac{\text{d}s(y;\bm{\eta}_{n})}{\text{d}y}\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}-\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))\right]\hskip 17.07182pt\text{if }c=1\\ -log\left[\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x})\right]\hskip 170.71652pt\text{if }c=0.\end{cases}$
@@ 3.4 Promoting fairness in soft survival trees
A crucial aspect in ML research is to deal with bias and fairness issues (Mehrabi et al., 2021). Indeed, ML models may perform an unfair discrimination by penalizing groups of data points characterized by sensitive attributes (features), like gender or ethnicity. Such discrimination often stems from inherent biases in the data, including the marginalization of historically disadvantaged groups (Pedreshi et al., 2008). To the best of our knowledge, no prior work accounts for fairness in survival trees.
@@
[…]
@@
[…]
@@ 4 Node-based decomposition algorithm for training soft survival trees
Decomposition algorithms have been devised to address similar challenges for other ML models such as Support Vector Machines (Chang and Lin, 2011), multilayer perceptrons (Grippo et al., 2015) and soft classification and regression trees (Amaldi et al., 2023; Consolo et al., 2025). The idea is to split the training problem into a sequence of smaller subproblems, where at each iteration the error function is optimized over a subset of variables (whose indices correspond to the so-called working set), while keeping the other variables fixed at their current values.
In this section we summarize the adaptation of the NOde-based DEComposition algorithm with Data points Reassignment heuristic (NODEC-DR) proposed in (Consolo et al., 2025) for training soft regression trees to soft survival trees, which will be referred to as NODEC-DR-SST. See C for a more detailed description.
@@
2:procedure NODEC-DR-SST($D,I,(\bm{\omega^{0}},\bm{\beta^{0}}),\varepsilon_{1},\varepsilon_{2},\varepsilon_{3},\theta$)
@@
4:     $error_{best}\leftarrow E(\bm{\hat{\omega}},\bm{\hat{\beta}})$
5:     while $it<M\_it\,and\,$ not termination criterion do
6:         for $s\in\{1,...,2^{D}-1\}$ $\triangleright$ loop over the branch nodes
@@
10:              $\bm{\omega}_{W^{k}_{B}}^{k+1}\leftarrow$UpdateBranchNode$(I,W^{k}_{B},\bm{\omega}^{k}_{W^{k}_{B}},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3})$triangleright update determined using heuristic
@@
In the LN Step, the new update $\bm{\beta}_{W^{k}_{L}}^{k+1}$ is determined by minimizing a negative log-likelihood function using any appropriate unconstrained optimization solver. This is in line with the standard approach in parametric and semi-parametric survival models, where the parameters of the survival functions are estimated via maximum likelihood. Similar to the BN Step, at each iteration $k$ the leaf node variables $\bm{\beta}$ are updated by considering the restricted training set $I_{s}$ of the branch node $s$, which is the root of the subtree defined by the working sets $W^{k}_{B}$ and $W^{k}_{L}$. It is worth pointing out that when using the objective function (4) to address fairness, the branch node variables $\bm{\omega}$ remain fixed at their current values during the LN Step. Thus, for any input vector $\mathbf{x}$, the unique leaf node $n_{\mathbf{x}}\in\tau_{L}$ in the corresponding HBP path can be readily determined.
@@ 5 Experimental results
After describing the considered datasets and the experimental settings, in Section 5.2 we compare the performance of SSTs trained via our node-based decomposition training algorithm with three widely used methods, namely, the survival tree Scikit-Survival (SkSurv) in (LeBlanc and Crowley, 1993), Conditional Inference Trees (CTree) in (Hothorn et al., 2006), and the CART-like Recursive Partitioning and Regression Trees (RPART) (Breiman et al., 1984). In Section 5.3, we illustrate how our SSTs enhance interpretability by inducing a partition of all the data points into clusters associated to the different leaf nodes with distinct survival functions. In Section 5.4, we show that the SST model and training formulation can be extended to promote group fairness in an unemployment dataset for a socioeconomic study (Romeo, 1999).
@@ 5.1 Datasets and experimental setting
In the numerical experiments, we consider 15 real-world survival datasets, all the 11 datasets from Zhang et al. (2024) and the first 4 listed in Table 1 of Huisman et al. (2024). Details about the type of application and the repositories containing the datasets are provided in A.1. Table 2 lists the names and the main characteristics of the datasets, namely, the number of features $p$, the number of data points $N$, and the censoring level expressed as a percentage. For all datasets, all input features are normalized to the $[0,1]$ range, while missing values are handled by imputing the mean for continuous features and the mode for categorical ones.
@@
[…]
@@
The experiments are carried out on a PC with Intel(R) Core(TM) i7-11370H CPU @ 3.30GHz with 16 GB of RAM. The SST model and the NODEC-DR-SST are implemented in Python 3.10.14. As previously mentioned, the performance of our SSTs is compared with that of the following implementation of the above-mentioned three methods: SkSurv (Pölsterl, 2020), CTree (Hothorn et al., 2015), and RPART (Therneau and Atkinson, 2019) [^4] The first from the Scikit-Survival package, and the last two from the CRAN package pec. Further information are reported in A.2.. Different measures are adopted to assess the discrimination and calibration capabilities, namely, Harrell’s C-index ($C_{H}$), Uno’s C-index ($C_{U}$), and the Cumulative Dynamic Area Under the Curve (CD-AUC) as discrimination measures, and the Integrated Brier Score (IBS) as calibration one. See B for a detailed description and references. Such performance measures are estimated by means of $k$-fold cross-validation, with $k=5.$ Due to the non convexity of the SST training formulation, 20 random initial solutions are considered for each fold. The final testing performance measures are then averaged over 5 folds and 20 initial solutions, that is, over 100 runs.
@@
As far as the initialization procedure for the Llog parametric distribution is concerned, we adopt the same clustering-based method proposed in (Consolo et al., 2025) for soft regression trees, based on the Kmeans function of the sklearn package. For SSTs with spline-based semiparametric survival functions, at each leaf node $n$ the initial solutions for the corresponding $\bm{\beta}_{n}$ vector are obtained using the method described in (Royston and Parmar, 2002). In particular, least-squares regression is applied to the logarithms of the cumulative hazards values (related to PH) or the cumulative odds values (related to PO), which are estimated from a random subset of uncensored data points (a fraction $\frac{1}{2^{D}}\%$ sampled with replacement) identified by fitting a standard Cox model. The COXHFitter function from the lifelines 0.29.0 package is used for fitting the Cox model. In the few cases (only a few runs for two datasets) where the COXHFitter function faces numerical issues, random values are assigned to the leaf node $\beta$ variables [^5] The values of the leaf node $\beta_{n}$ parameters are selected to guarantee the positivity of the derivative $\frac{\text{d}s(y;\bm{\eta})}{\text{d}y}$, which is the argument of the logarithm in the negative log-likelihood..
@@
Concerning the spline-based PH and PO survival functions at the SSTs leaf nodes, we consider $m=2$ internal knots, placed according to Table 1 in (Royston and Parmar, 2002).
@@
As done in (Consolo et al., 2025), we adopt an early-stopping strategy for the NODEC-DR algorithm, where $M_{it}=10$, $k_{0}>M_{it}$, and the hyperparameters $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$, $\varepsilon_{3}=0.4$, and $\zeta=0.8$ are used for the data points reassignment heuristic. To improve numerical stability, an $\ell_{2}$-penalty term is added to the error function (2). This term amounts to $\ell_{2}$ norm of the survival function variables $\bm{\beta}_{n}$ in the leaf nodes, multiplied by a constant value of $2$.

### v3

@@ 3.3 Parametric and spline-based semiparametric models for survival functions
In statistical modeling, several alternatives have been proposed for the estimation of the survival and hazard functions. Traditionally, medical researchers and biostatisticians have relied on Cox regression (cox1972regression) to handle censored survival data. This model assumes that the hazard rate for an event is a linear combination of the features, with the hazard function for an input vector $\mathbf{x}$ expressed as:
@@
In the Cox model, the choice of $h_{0}(t)$ is critical since its estimation in high-dimension is prone to overfitting and may complicate the validation (gelfand2000proportional; royston2002flexible). Considering more sophisticated hazard functions can lead to more accurate estimates and offer deeper insights into the phenomenon under investigation (e.g., capturing the progression of a disease over time in medical studies). Moreover, the classical proportional hazards assumption of the Cox model may not be suitable for all contexts [^2] Although the Cox model allows for non-proportional hazard assumption, such as through time-varying regression coefficients, there is no widely accepted or intuitive method for implementing this adjustment..
To overcome these limitations, more flexible statistical approaches, including parametric distribution and spline-based semiparametric ones, have been proposed. These approaches are well-suited to capture complex patterns in the data, such as time-varying hazards, interval censoring, frailties, and multiple responses across diverse datasets or time scales (aalen2008survival).
@@ 3.3.1 Parametric distribution models
All parameters may be linearly dependent on the features through a transformation $g$ such that $g(\alpha_{r}(\mathbf{x}))=\gamma_{0}+\bm{\gamma}^{T}\mathbf{x}$ for $r=1,\dots,R$. Typically, $g$ is the logarithm function for positive parameters, and the identity function for parameters which are unrestricted in sign. For a review of the parametric distributions, the reader is referred to (lawless2014parametric; klein2014handbook).
In this work, we assume that only the primary parameter $\mu$ linearly depends on the features (i.e., $\text{log}(\mu(\mathbf{x}))=\gamma_{0}+\bm{\gamma}^{T}\mathbf{x}$). Thus, for each leaf node $n\in\tau_{L}$, the vector $\bm{\beta}_{n}$ contains $p+1+m$ components. The first $p+1$ components correspond to the number of variables related to $\mu$, and the m=R1m=R-1 remaining ones to the ancillary parameters.
@@ 3.3.2 Spline-based semiparametric models
For the estimation of the semiparametric survival functions, we adopt the methodology proposed by royston2002flexible, which assumes flexible spline-based models to take into account proportional hazards and proportional odds. Other options (see, e.g., younes1997link; shen1998propotional; royston2011use; bremhorst2016flexible) that preserve the smoothness of formulation (7) can also be considered.
Given any input vector $\mathbf{x}$, royston2002flexible assume that the transformation of the survival function $g(S_{\mathbf{x}}(t;\bm{\beta}))$ is equal to the transformation of a baseline survival function $S_{0}(t)$ plus a linear combination of the features:
$g(S_{\mathbf{x}}(t;\bm{\beta}))=g(S_{0}(t))+\bm{\gamma}^{T}\mathbf{x},$ (8)
@@
Concerning the transformation $g$, following (aranda1981two), royston2002flexible define:
$g(x;\theta)=\text{log}\frac{x^{-\theta}-1}{\theta},$
where the value of $\theta$ determines whether the proportional odds or the proportional hazards model is obtained. When $\theta=1$, we have the proportional odds model (i.e., the classical logistic model), while as $\theta$ tends to 0 we recover the proportional hazards model.
Following royston2002flexible, we estimate the baseline survival function $S_{0}(t)$ using spline-based models. Specifically, in the proportional odds case we model the logarithm of the baseline cumulative odds $\text{log}\left(\tfrac{1-S_{0}(t)}{S_{0}(t)}\right)$, while in the proportional hazards case we model the logarithm of the baseline hazard $\text{log}H_{0}(t)$, both as natural cubic splines of log-time $y=\text{log}(t)$.
From (8), this leads to the Proportional Odds (PO) spline model:
@@
The complexity of such splines is determined by the number of degrees of freedom, which is equal to $m+1$. For the placement of internal knots, royston2002flexible propose to use the percentiles of the distribution of uncensored log survival times. Further details can be found in Table 1 of royston2002flexible.
According to (6), for any leaf node $n\in\tau_{L}$ and any right-censored data point $(\textbf{x},t,c)$, the corresponding negative log-likelihood $L^{-}_{n}$ in the PO model is:
$L^{-}_{n}({\mathbf{x}},t,c;\bm{\beta}_{n})=\begin{cases}-log\left[\frac{1}{t}\frac{\text{d}s(y;\bm{\eta}_{n})}{\text{d}y}\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x})(1+\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))^{-2}\right]\hskip 0.0pt\text{if }c=1\\ -log\left[(1+\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))^{-1}\right]\hskip 128.0374pt\text{if }c=0,\end{cases}$ (9)
whereas in the PH model it is as follows:
$L^{-}_{n}({\mathbf{x}},t,c;\bm{\beta}_{n})=\begin{cases}-log\left[\frac{1}{t}\frac{\text{d}s(y;\bm{\eta}_{n})}{\text{d}y}\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}-\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x}))\right]\hskip 27.03003pt\text{if }c=1\\ -log\left[\text{exp}(s(y;\bm{\eta}_{n})+\bm{\gamma}_{n}^{T}\mathbf{x})\right]\hskip 167.87108pt\text{if }c=0.\end{cases}$ (10)
@@ 3.4 Promoting fairness in soft survival trees
A crucial aspect in ML research is to deal with bias and fairness issues (mehrabi2021survey). Indeed, ML models may perform an unfair discrimination by penalizing groups of data points characterized by sensitive attributes (features), like gender or ethnicity. Such discrimination often stems from inherent biases in the data, including the marginalization of historically disadvantaged groups (pedreshi2008discrimination). Even if a sensitive feature is removed or ignored, individuals in the sensitive group may still be disadvantaged, since other features may often be strongly correlated with the group membership.
During the past decade, several attempts have been made to train decision trees while accounting for fairness issues. Recent contributions include MILP-based methods to incorporate group fairness into decision trees, combining feature selection with fairness constraints (aghaei2019learning; wang2022synthesizing; jo2023learning), as well as a dynamic programming approach that builds classification trees satisfying group fairness requirements (van2022fair). To the best of our knowledge, no prior work takes into account fairness in survival trees.
@@
$\sum_{i\in G}\sum_{j\in\bar{G}}\int^{t_{\text{max}}}_{t_{\text{min}}}(S_{\textbf{x}_{i}}(t;\bm{\beta}_{n_{\mathbf{x}_{i}}})-S_{\textbf{x}_{j}}(t;\bm{\beta}_{n_{\mathbf{x}_{j}}}))^{2}dt,$ (11)
@@
$\begin{split}\min_{\bm{\omega},\bm{\beta}}\frac{1}{N}\sum_{i\in I}\sum_{k\in\tau_{L}}P_{in}(\bm{\omega})L^{-}_{n}(\mathbf{x}_{i},t_{i},c_{i};\bm{\beta}_{n})+\rho\sum_{i\in G}\sum_{j\in\bar{G}}\int^{t_{\text{max}}}_{t_{\text{min}}}(S_{\textbf{x}_{i}}(t;\bm{\beta}_{n_{\mathbf{x}_{i}}})-S_{\textbf{x}_{j}}(t;\bm{\beta}_{n_{\mathbf{x}_{j}}}))^{2}dt,\end{split}$ (12)
@@ 4 Node-based decomposition algorithm for training soft survival trees
Decomposition algorithms have been devised to address similar challenges for other ML models such as Support Vector Machines (chang2011libsvm), multilayer perceptrons (grippo2015decomposition) and soft classification and regression trees (amaldi2021multivariate; consolo2025). The idea is to split the training problem into a sequence of smaller subproblems, where at each iteration the error function is optimized over a subset of variables (whose indices correspond to the so-called working set), while keeping the other variables fixed at their current values.
In this section we summarize the adaptation of the NOde-based DEComposition algorithm with Data points Reassignment heuristic (NODEC-DR) proposed in (consolo2025) for training soft regression trees to soft survival trees, which will be referred to as NODEC-DR-SST. See C for a more detailed description.
@@
[…]
@@
[…]
5:  while it<M_itandit<M\_it\,and\, not termination criterion do
@@
[…]
10:     $\bm{\omega}_{W^{k}_{B}}^{k+1}\leftarrowUpdateBranchNode$$(I,W^{k}_{B},\bm{\omega}^{k}_{W^{k}_{B}},\varepsilon_{1},\varepsilon_{2},\varepsilon_{3})$ $\triangleright$ update determined using heuristic
@@
In the LN Step, the new update $\bm{\beta}_{W^{k}_{L}}^{k+1}$ is determined by minimizing a negative log-likelihood function using any appropriate unconstrained optimization solver. This is in line with the standard approach in parametric and semiparametric survival models, where the parameters of the survival functions are estimated via maximum likelihood. Similar to the BN Step, at each iteration $k$ the leaf node variables $\bm{\beta}$ are updated by considering the restricted training set $I_{s}$ of the branch node $s$, which is the root of the subtree defined by the working sets $W^{k}_{B}$ and $W^{k}_{L}$. It is worth pointing out that when using the objective function (12) to address fairness, the branch node variables $\bm{\omega}$ remain fixed at their current values during the LN Step. Thus, for any input vector $\mathbf{x}$, the unique leaf node $n_{\mathbf{x}}\in\tau_{L}$ in the corresponding HBP path can be readily determined.
@@ 5 Experimental results
After describing the considered datasets and the experimental settings, Section 5.2 provides a brief overview of the performance measures adopted in this work. In Section 5.3 we compare the performance of SSTs trained via our node-based decomposition training algorithm with three widely used methods, namely, the survival tree Scikit-Survival (SkSurv) in (leblanc1993survival), Conditional Inference Trees (CTree) in (hothorn2006unbiased), and the CART-like Recursive Partitioning and Regression Trees (RPART) (breimanclassification). In Section 5.4, we illustrate how our SSTs enhance interpretability by inducing a partition of all the data points into clusters associated to the different leaf nodes with distinct survival functions. In Section 5.5, we show that the SST model and training formulation can be extended to promote group fairness in an unemployment dataset for a socioeconomic study (romeo1999conducting).
@@ 5.1 Datasets and experimental setting
In the numerical experiments, we consider 15 real-world survival datasets, all the 11 datasets from zhang2024optimal and the first 4 listed in Table 1 of huisman2024optimal. Details about the type of application and the repositories containing the datasets are provided in A.1. Table 2 lists the names and the main characteristics of the datasets, namely, the number of features $p$, the number of data points $N$, and the censoring level expressed as a percentage. For all datasets, all input features are normalized to the $[0,1]$ range, while missing values are handled by imputing the mean for continuous features and the mode for categorical ones.
@@
Aids 1151 11 91.7 $\%$
Aids_death 1151 11 91.7 $\%$
Aids2 2839 4 38.0 $\%$
Churn 2000 12 53.4 $\%$
Credit 1000 19 30.0 $\%$
Dialysis 6805 4 76.4 $\%$
Employee 11992333The initial dataset contained 15000 data points, but we applied the preprocessing steps outlined in fotso2019pysurvival, as described in https://square.github.io/pysurvival/tutorials/employee_retention.html. 7 77.2 $\%$
Flchain 7478 9 72.5 $\%$
Framingham 4658 7 68.5 $\%$
Gbsg2 686 8 56.4 $\%$
Maintenance 1000 5 60.3 $\%$
Uissurv 628 12 19.1 $\%$
Unempdur 3241 6 38.7 $\%$
Veterans 137 6 6.6 $\%$
Whas500 500 14 57.0 $\%$
@@
The experiments are carried out on a PC with Intel(R) Core(TM) i7-11370H CPU @ 3.30GHz with 16 GB of RAM. The SST model and the NODEC-DR-SST are implemented in Python 3.10.14. As previously mentioned, the performance of our SSTs is compared with that of the following implementation of the above-mentioned three methods: SkSurv (polsterl2020scikit), CTree (hothorn2015ctree), and RPART (therneau2021rpart) [^4] The first from the Scikit-Survival package, and the last two from the CRAN package pec. Further information are reported in A.2.. Different measures are adopted to assess the discrimination and calibration capabilities, namely, Harrell’s C-index ($C_{H}$), Uno’s C-index ($C_{U}$), and the Cumulative Dynamic Area Under the Curve (CD-AUC) as discrimination measures, and the Integrated Brier Score (IBS) as calibration one. In Section 5.2, we provide a concise overview of the four measures considered in this study (a detailed description and references can be found in B), with particular emphasis on $C_{H}$ and IBS. These two measures, which assess discrimination and calibration respectively, are the ones whose results are reported in Section 5.3 (results for $C_{U}$ and CD-AUC on the same experiments are provided in D). Such performance measures are estimated by means of $k$-fold cross-validation, with $k=5.$ Due to the non convexity of the SST training formulation, 20 random initial solutions are considered for each fold. The final testing performance measures are then averaged over 5 folds and 20 initial solutions, that is, over 100 runs. Furthermore, to evaluate the statistical significance of the comparison between the best parametric and semiparametric SST and the best benchmark model, we perform statistical tests on the mean differences of the testing accuracies for each of the four considered measures. In particular, we employ the one-sided non-parametric Wilcoxon signed-rank test, which is preferred over the classical t-test to avoid the restrictive assumption of normally distributed differences.
@@
As far as the initialization procedure for the Llog parametric distribution is concerned, we adopt the same clustering-based method proposed in (consolo2025) for soft regression trees, based on the Kmeans function of the sklearn package. For SSTs with spline-based semiparametric survival functions, at each leaf node $n$ the initial solutions for the corresponding $\bm{\beta}_{n}$ vector are obtained using the method described in (royston2002flexible). In particular, least-squares regression is applied to the logarithms of the cumulative hazards values (related to PH) or the cumulative odds values (related to PO), which are estimated from a random subset of uncensored data points (a fraction $\frac{1}{2^{D}}\%$ sampled with replacement) identified by fitting a standard Cox model. The COXHFitter function from the lifelines 0.29.0 package is used for fitting the Cox model. In the few cases (only a few runs for two datasets) where the COXHFitter function faces numerical issues, random values are assigned to the leaf node $\beta$ variables [^5] The values of the leaf node $\beta_{n}$ parameters are selected to guarantee the positivity of the derivative $\frac{\text{d}s(y;\bm{\eta})}{\text{d}y}$, which is the argument of the logarithm in the negative log-likelihood..
@@
Concerning the spline-based PH and PO survival functions at the SSTs leaf nodes, we consider $m=2$ internal knots, placed according to Table 1 in (royston2002flexible).
@@
As done in (consolo2025), we adopt an early-stopping strategy for the NODEC-DR algorithm, where $M_{it}=10$, $k_{0}>M_{it}$, and the hyperparameters $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.3$, $\varepsilon_{3}=0.4$, and $\zeta=0.8$ are used for the data points reassignment heuristic. To improve numerical stability, an $\ell_{2}$-penalty term is added to the error function (7). This term amounts to $\ell_{2}$ norm of the survival function variables $\bm{\beta}_{n}$ in the leaf nodes, multiplied by a constant value of $2$.
5.2 Discrimination and calibration measures
Performance measures for survival models are typically concerned with two aspects: discrimination and calibration. Discrimination evaluates a model’s ability to separate individuals who experience the event from those who do not, by assigning higher predicted probabilities to the former ones. Calibration assesses the agreement between predicted probabilities and observed outcomes.
Among the discrimination measures, Harrell’s Concordance Statistic ($C_{H}$) is the most widely applied in survival analysis (yan2008investigating). $C_{H}$ is based on the principle that a reliable survival model should assign lower survival probabilities to individuals with shorter times-to-event. To define the concordance measure, it is necessary to introduce the notions of comparable and concordant pairs of individuals. A pair $(i,j)$ of individuals is comparable if the order of the events can be determined and it is concordant when the individual with the earlier event receives a lower predicted survival probability. Since survival predictions depend on time, concordance is assessed at $\min\{t_{i},t_{j}\}$. Formally, Harrell’s C-index is defined as:
$C_{H}=\frac{\sum\limits_{i}\sum\limits_{j}c_{i}\cdot\mathds{1}_{t_{i}<t_{j}}\cdot\left(\mathds{1}_{\hat{S}_{\mathbf{x}_{i}}(t_{i})<\hat{S}_{\mathbf{x}_{j}}(t_{i})}+0.5\cdot\mathds{1}_{\hat{S}_{\mathbf{x}_{i}}(t_{i})=\hat{S}_{\mathbf{x}_{j}}(t_{i})}\right)}{\sum\limits_{i}\sum\limits_{j}c_{i}\cdot\mathds{1}_{t_{i}<t_{j}}}.$
$C_{H}$ ranges from 0 to 1 and higher values indicate better model discrimination.
Since $C_{H}$ may be biased for datasets with a high proportion of censored data points, uno2011c introduced an extended version of the concordance index ($C_{U}$) based on Inverse Probability of Censoring Weights (see B for details), which which is more robust under extensive censoring. Like $C_{H}$, $C_{U}$ ranges from 0 to 1, with higher values indicating better discrimination.
Another popular discrimination measure used in survival analysis is the Cumulative Dynamic Area Under the Curve (CD-AUC). This measure extends the classical AUC for binary classification to time-to-event data by using time-dependent Receiver Operating Characteristic functions (steyerberg2009clinical), where at each time $t$ true positives are individuals experiencing the event at or before $t$, and true negatives are those remaining event-free beyond $t$. The CD-AUC summarizes time-dependent AUC values over a chosen interval, yielding an overall measure of discrimination across the the considered time interval. Like the standard AUC, its values range from 0 to 1, with higher values indicating better discrimination. The reader is referred to B for further details.
As to calibration, a widely used measure is the Brier Score (BS), originally developed for probabilistic classification and later adapted to censored survival data (graf1999assessment). In survival analysis, $BS(t)$ at time $t$ is the mean squared error between the observed outcomes and the predicted survival probabilities $\hat{S}_{\mathbf{x}_{i}}(t)$, adjusted through inverse probability of censoring weights:
$\color[rgb]{0.0,0.0,0.0}BS(t)=\frac{1}{N}\sum_{i=1}^{N}\left(\frac{(\hat{S}_{\mathbf{x}_{i}}(t)-0)^{2}}{\hat{G}(t_{i})}\cdot\mathds{1}_{t_{i}\leq t,c_{i}=1}+\frac{(\hat{S}_{\mathbf{x}_{i}}(t)-1)^{2}}{\hat{G}(t)}\cdot\mathds{1}_{t_{i}>t}\right),$}
where $\hat{G}$ represents the Kaplan–Meier estimate of the censoring distribution $\mathbf{c}$.
Similar to CD-AUC, the Integrated Brier Score (IBS) defined as
$\color[rgb]{0.0,0.0,0.0}\text{IBS}=\frac{1}{t_{\text{max}}}\int_{0}^{t_{\text{max}}}BS(t)\,dt$}
where $t_{\text{max}}=\max_{i=1}^{N}t_{i}$ being the latest observed time point, provides a global assessment of the odel performance over the time interval $0,t_{\text{max}}]$. BS and IBS range from 0 to 1, with lower values corresponding to better calibration, see B for further details.